A car traveling at constant velocity of 20m/s passes a stationary police car and 15 seconds later the police car accelerates uniformly at 4m/s to reach a speed of 30m/s with which it continues. a) draw a velocity time graph. b) find where the second car overtakes the first car. c) find the distance up to which the second car overtakes the first car


1
Expert's answer
2022-03-13T18:46:41-0400

Solution.

"v_c=20m\/s;"

"t_c=15s;"

"v_{p0}=0m\/s;"

"v_p=30m\/s;"

"a_p=4m\/s;"

"a)"


"b) S_c=v_c\sdot t_c=20\sdot15=300m;"

"a_p=\dfrac{v_p-v_{p0}}{t_p}\implies t_p=\dfrac{v_p-v_{p0}}{a_p};"

"t_p=\dfrac{30-0}{4}=7.5s;"

"x_{c0}=S_c+v_ct_p=300+20\sdot7.5=450m;"

"x_c=450+20t;"

"x_p=30t;"

"x_c=x_p;"

"450+20t=30t;"

"10t=450;"

"t=45s;"

"c)x_c=450+20t=450+20\sdot45=1350m;"

Answer: "b)t=45s;"

"c) x_c=1350m."